# Nesterov acceleration in benignly nonconvex landscapes

### v1

@@ 3.1. Optimization in continuous time
(3) $\ddot{x}+\gamma\,\dot{x}=-\nabla f(x),\qquad x(0)=x_{0},\qquad\dot{x}(0)=0.$
@@
[Continuous time convergence guarantee] Assume that $f$ satisfies the assumptions of Section 2.1, $\gamma=2\sqrt{\mu}$ and that $x_{0}$ satisfies $f(x_{0})-\inf f<\alpha$ where $\alpha$ is as in Section 2.1. Then there exists a unique solution $x(t)=x_{t}$ of (3) and $x_{t}\in\mathcal{U}_{\alpha}$ for all $t>0$. Furthermore, $x$ is $C^{2}$-smooth and
@@
The proof can be found in Appendix B. There, we prove that the function
@@
satisfies ${\mathcal{L}}(t)\leq e^{-\sqrt{\mu}t}{\mathcal{L}}(0)$. The function ${\mathcal{L}}$ can be considered as a modified energy with potential energy $f-\inf f+\frac{\mu}{2}\|x-\pi(x)\|^{2}$ and kinetic energy $\frac{1}{2}\|\dot{x}\|^{2}$, but with a more complex quadratic term in which velocity and position interact. Going back to (Su et al., 2016), this strategy of proof is common in convex optimization. Our contribution is the combination with Lemma 1 to control additional terms in the more general geometric setting.
@@
The rate of decay is ‘accelerated’ since the rate of decay $e^{-\sqrt{\mu}\,t}$ depends on the square root of the surrogate of the convexity parameter rather than $\mu$. For comparison, under the same assumptions, the gradient flow would decay like $e^{-\mu t}$ (since first order strong convexity with respect to the closest minimiser implies the PL condition). We see more clearly why we speak of acceleration in discrete time (see the discussion below Theorem 11).
@@
3.2. Deterministic optimization in discrete time
@@
(4) $\left\{\begin{array}[]{rll}\dot{x}&=v\\ \dot{v}&=-2\sqrt{\mu}\,v-\nabla f(x),\end{array}\right.\qquad x(0)=x_{0},\qquad v(0)=0.$
@@
(5) $x^{\prime}_{n}=x_{n}+\sqrt{\eta}v_{n},\qquad g_{n}=\nabla f(x^{\prime}_{n}),\qquad x_{n+1}=x^{\prime}_{n}-\eta g_{n},\qquad v_{n+1}=\rho\big{(}v_{n}-\sqrt{\eta}g_{n}\big{)}$
@@
i.e. the scheme is a time discretization of the heavy ball system (4) with time-step size $\sqrt{\eta}$. The square root is chosen for consistency with the literature. This scheme is a geometrically intuitive reparametrization of Nesterov’s accelerated gradient descent algorithm, except for the fact that Nesterov’s scheme typically begins with the gradient descent step rather than the momentum step.
In discrete time, we need to make additional quantitative regularity assumptions on both $f$ and the projection map $\pi$ in order to ensure that the time step size is sufficiently small to recover the continuous time behavior. Note that the Hessian of the objective function $f$ is non-negative semi-definite on the set of global minimizers $\mathcal{M}$, i.e. it stands to reason that any negative eigenvalues of $D^{2}f$ should be small, at least close to $\mathcal{M}$. We note the following.
@@
(6) $\langle\nabla f(x),x-\pi(x)\rangle\geq f(x)-f(\pi(x))+\frac{\mu}{2}\,\|x-\pi(x)\|^{2}.$
@@
(7) $\langle\nabla f(x+v),v\rangle\geq f(x+v)-f(x)-\frac{\varepsilon}{2}\,\|v\|^{2}$
@@
The proof is a more complicated version of that of Theorem 7 in which we control for difficulties stemming from the time-discrete setting. It can be found in Appendix B.2.
In Theorem 11, we see more clearly than in Theorem 7 why we talk of acceleration: While gradient descent would achieve a decay rate of $(1-\mu/L)^{n}$ with the commonly proposed step size $\eta=1/L$ in discrete time based on our assumptions, Nesterov’s method achieves decay like $(1-\sqrt{\mu/L})^{n}$ with $\eta=1/L$. If $\mu/L$ is close to zero, Nesterov’s method converges much faster than gradient descent.
@@
However, if the eigenvalues of $D^{2}f$ are as negative as $-L$, the step size $\eta=\mu/L^{2}$ is so small that it does not improve upon gradient descent with step size $\eta=1/L$ since $\sqrt{\mu\eta}=\mu/L$ in this case. Thus, if $f$ is too far from being convex, acceleration may not be achievable in discrete time. Notably, especially close to the set of global minimizers, we can picture $\varepsilon$ as a small compared to $L$.
The condition on $\varepsilon$ is less restrictive in Theorem 15 for the stochastic setting: As we contend with the major challenge of stochastic gradient estimates, the best achievable rate of convergence becomes much slower. In this setting, the additional complication of tangential drift becomes less dire.
3.3. Stochastic optimization in discrete time
In typical applications in deep learning, the gradient $\nabla f$ of the objective function/loss function $f$ is prohibitively expensive to evaluate, but we have access to stochastic estimates of the true gradient. In this section, in addition to the assumptions of Theorem 11, we assume that we are given a probability space $(\Omega,\mathcal{A},\mathbb{Q})$ and a (measurable) function $g:\mathbb{R}^{d}\times\Omega\to\mathbb{R}^{d}$ such that
(8) ${\mathbb{E}}_{\omega\sim\mathbb{Q}}\big{[}g(x,\omega)\big{]}=\nabla f(x),\qquad{\mathbb{E}}_{\omega\sim\mathbb{Q}}\big{[}\|g(x,\omega)\|^{2}\big{]}<+\infty\qquad\forall\ x\in\mathbb{R}^{d}.$
@@ 3.3. Stochastic optimization in discrete time
(9) ${\mathbb{E}}_{\omega\sim\mathbb{Q}}\big{[}\|g(x,\omega)-\nabla f(x)\|^{2}\big{]}\leq\sigma_{a}^{2}+\sigma_{m}^{2}\|\nabla f(x)\|^{2}\qquad\forall\ x\in\mathbb{R}^{d}.$
@@
Let us consider the purely additive case first. We follow the scheme (5), but we replace the deterministic gradient $\nabla f(x_{n}^{\prime})$ by $g_{n}=g(x_{n}^{\prime},\omega_{n})$ where $\omega_{0},\omega_{1},\dots$ are drawn from $\Omega$ independently of each other and the initial condition $x_{0}$ with law $\mathbb{Q}$. This framework allows e.g. for minibatch sampling (but assumes that all batches are drawn independently from the dataset without correlation between the batch in one timestep and the next).
@@
Thus Nesterov’s method reduces $f(x_{n})$ below a ‘noise level’ proportional to $\sigma_{a}^{2}\sqrt{\eta}/\sqrt{\mu}$ at a linear rate $(1-\sqrt{\mu\eta})^{n}$ in the presence of additive noise. The analogous bound for stochastic gradient descent was obtained (in the more general setting of PL functions) in (Karimi et al., 2016, Theorem 4) as
@@
With a fixed positive learning rate $\eta$, generally $f(x_{n})\not\to 0$. We therefore consider a sequence of decreasing step sizes.
@@
(10) $x_{n}^{\prime}=x_{n}+\sqrt{\alpha}v_{n},\qquad x_{n+1}=x^{\prime}_{n}-\eta g_{n},\qquad v_{n+1}=\rho_{n}\big{(}v_{n}-\sqrt{\alpha}g_{n})$
@@
[Additive and multiplicative noise] Assume that $f,P,x^{*}$ are as in Theorem 11 and that $g$ is a family of gradient estimators such that (8) and (9) hold for some $\sigma_{a},\sigma_{m}\geq 0$.
Assume that the sequences $x_{n},x_{n}^{\prime},v_{n}$ are generated by the AGNES scheme (10) with parameters
$0<\eta\leq\frac{1}{L(1+\sigma^{2})},\qquad\rho=\frac{1-\sqrt{\frac{\mu\eta}{1+\sigma_{m}^{2}}}}{1+\sqrt{\frac{\mu\eta}{1+\sigma_{m}^{2}}}},\qquad\alpha=\frac{1-\sqrt{\mu(1+\sigma_{m}^{2})\eta}}{1-\sqrt{\mu(1+\sigma_{m}^{2})\eta}+\sigma_{m}^{2}}\,\eta.$
@@
While at a glance it appears that the multiplicative noise is helping us reduce the additive error term, this is merely a consequence of the small learning rate which it forced upon us.
The proof of Theorem 15 is given in Appendix B.4. As indicated in Remark 12, the condition on the negative eigenvalues is relaxed to $\varepsilon\leq\sqrt{L\mu}(1+\sigma_{m}^{2})$ with the largest admissible step size $\eta=1/(L(1+\sigma_{m}^{2}))$ as the issues stemming from tangential drift pale in comparison to those stemming from stochastic gradient estimates. For comparison, if we chose the same $\eta$ in Theorem 11, we could only allow for $\varepsilon\leq\sqrt{\mu L}\,\sqrt{1+\sigma_{m}^{2}}<\varepsilon\sqrt{\mu L}(1+\sigma_{m}^{2})$, but we would obtain a rate of convergence of $1-\sqrt{\mu/{L(1+\sigma_{m}^{2})}}$. In the stochastic case, we only achieve $1-\sqrt{\mu/L}/(1+\sigma_{m}^{2})$. Thus the limiting factor is the stochastic noise, not the geometry of $f$.
@@
4. Conclusion
Figure 4. We compare the trajectories of gradient descent and Nesterov’s algorithm for the objective function $f$ in Example 2 with $R=6$ and $\varepsilon=0.075$ (left), $\varepsilon=0.08$ (middle) and $\varepsilon=0.085$ (right). Evidently, if $\varepsilon\sqrt{1+4R^{2}}$ is very close to the threshold value 1, gradient descent outperforms Nesterov’s algorithm with the theoretically guaranteed parameters.
We have proved that first order momentum-based methods accelerate convergence in a much more general setting than convex optimization with many geometric features motivated by realistic loss landscapes in deep learning. The models we studied include the heavy ball ODE and deterministic and stochastic optimization schemes in discrete time under various noise assumptions.
Let us conclude by discussing the guarantees obtained in this work. In the case of quadratic objective functions or functions modelled on them (such as the distance function from a manifold), the PL constant is essentially the same as the parameter of strong convexity. For the function $f$ in Example 2 on the other hand, the PL constant is noticeable larger than the parameter of strong convexity with respect to the unique minimizer. In Figure 4, we observe that with the theoretically guaranteed parameters $\eta=1/L$ and $\rho=(1-\sqrt{\mu\eta})/(1+\sqrt{\mu\eta})$, gradient descent may converge faster in some cases.
Appendix A Proofs of Lemmas 1 and 6: Geometry of the energy landscape
@@
is convex. In particular, the derivative matrix $D\pi=D^{2}\xi$ is symmetric and non-negative semi-definite and thus
@@
Step 2. The geometry of $f$. Since $\mathcal{M}$ is the set of minimizers of $f$, the Hessian $D^{2}f(z)$ is non-negative semi-definite for every $z\in\mathcal{M}$. As $f$ is constant on $\mathcal{M}$, we for every curve $\gamma:(-\varepsilon,\varepsilon)\to\mathcal{M}$ we have
@@ Appendix B Proofs of Acceleration in Optimization
B.1. Proof of Theorems 7 and 8: Optimization in continuous time
@@
Next, using the $\mu$-strong convexity of $f$ with respect to the closest minimizer, we have
@@
Substituting this into the bound on $L^{\prime}(t)$ and simplifying gives:
@@
B.2. Proof of Theorem 11: Acceleration in Discrete Time (Deterministic setting)

### v2

@@ 3.1 Optimization in continuous time
$\ddot{x}+\gamma\,\dot{x}=-\nabla f(x),\qquad x(0)=x_{0},\qquad\dot{x}(0)=0.$ (3)
@@
[Continuous time convergence guarantee] Assume that $f$ satisfies the assumptions of Section 2.1, $\gamma=2\sqrt{\mu}$ and that $x_{0}$ satisfies $f(x_{0})<\alpha$ where $\alpha$ is as in Section 2.1. Then there exists a unique solution $x(t)=x_{t}$ of (3) and $f(x_{t})<\alpha$ for all $t>0$. Furthermore, $x$ is $C^{2}$-smooth and
@@
The proof can be found in Appendix B. There, we prove that the Lyapunov function
@@
satisfies ${\mathcal{L}}(t)\leq e^{-\sqrt{\mu}t}{\mathcal{L}}(0)$. The function ${\mathcal{L}}$ can be considered as a modified energy with potential energy $f-\inf f+\frac{\mu}{2}\|x-\pi(x)\|^{2}$ and kinetic energy $\frac{1}{2}\|\dot{x}\|^{2}$, but with a more complex quadratic term in which velocity and position interact. Going back to (Su et al., 2016), this strategy of proof is common in convex optimization. The main difficulty in our more general geometric setting is that the time derivative of $\pi(x_{t})$ (i.e. tangential velocity) is not zero in general, unless the minimizer is unique in $\mathcal{U}_{\alpha}.$ Thus, we have to additionally control for the interaction of the tangential velocity with the other terms in the Lyapunov sequence. One key observation that helps us is that the connecting line $x_{t}-\pi(x_{t})$ is orthogonal to the velocity of $\pi(x_{t})$. The other technical tool is the following lemma.
Lemma 7.
Let $\mathcal{M}$ be a $C^{2}$-submanifold of $\mathbb{R}^{d}$ and $\mathcal{U}$ an open set containing $\mathcal{M}$ such that there exists a unique closest point projection $\pi:\mathcal{U}\to\mathcal{M}$. Let $x:(-\varepsilon,\varepsilon)\to\mathcal{U}$ be a $C^{1}$-curve and $z(t):=\pi\circ x(t)$. Then $\langle\dot{x},\dot{z}\rangle\geq 0$ on $(-\varepsilon,\varepsilon)$.
@@
The rate of decay here is $e^{-\sqrt{\mu}\,t}$, as compared to $e^{-\mu t}$ for gradient flow. We see more clearly why we speak of acceleration in discrete time (see the discussion below Theorem 11).
@@
3.2 Deterministic optimization in discrete time
@@
$\left\{\begin{array}[]{rll}\dot{x}&=v\\ \dot{v}&=-2\sqrt{\mu}\,v-\nabla f(x),\end{array}\right.\qquad x(0)=x_{0},\qquad v(0)=0.$ (4)
@@
$x^{\prime}_{n}=x_{n}+\sqrt{\eta}v_{n},\qquad g_{n}=\nabla f(x^{\prime}_{n}),\qquad x_{n+1}=x^{\prime}_{n}-\eta g_{n},\qquad v_{n+1}=\rho\big{(}v_{n}-\sqrt{\eta}g_{n}\big{)}$ (5)
@@
i.e. the scheme is a time discretization of the heavy ball system (4) with time-step size $\sqrt{\eta}$. The square root is chosen for consistency with the literature. This scheme is a geometrically intuitive reparametrization of Nesterov’s accelerated gradient descent algorithm, except for the fact that Nesterov’s scheme typically begins with the gradient descent step rather than the momentum step (see Gupta et al., 2024, Appendix B, for a proof of equivalence).
In discrete time, we need to make additional quantitative regularity assumptions on both $f$ and the projection map $\pi$ in order to ensure that the time step size is sufficiently small to recover the continuous time behavior. Note that the Hessian of the objective function $f$ is positive semi-definite on the set of global minimizers $\mathcal{M}$, i.e. it stands to reason that any negative eigenvalues of $D^{2}f$ should be small, at least close to $\mathcal{M}$. We note the following.
@@
$\langle\nabla f(x),x-\pi(x)\rangle\geq f(x)-f(\pi(x))+\frac{\mu}{2}\,\|x-\pi(x)\|^{2}.$ (6)
@@
$\langle\nabla f(x+v),v\rangle\geq f(x+v)-f(x)-\frac{\varepsilon}{2}\,\|v\|^{2}$ (7)
@@
The proof, given in Appendix B.2, builds on similar ideas as Theorem 6, with additional complications introduced by the discrete-time setting. We have to modify the usual Lyapunov sequence used for strongly convex functions. We treat the tangential and normal components of the velocity (i.e. $Pv_{n}$ and $P^{\bot}v_{n}=(I-P)v_{n}$) separately, and carefully choose coefficients to ensure that the following Lyapunov sequence decays at each step,
${\mathcal{L}}_{n}=f(x_{n})-\inf f+\frac{1}{2}\|\ P^{\bot}v_{n}+\sqrt{\mu}(x_{n}^{\prime}-\pi(x_{n}^{\prime}))\|^{2}+\frac{(1+\sqrt{\mu\eta})^{2}}{2(1-\sqrt{\mu\eta})}\|Pv_{n}\|^{2}.$
In Theorem 11, we see more clearly than in Theorem 6 why we talk of acceleration: While gradient descent would achieve a decay rate of $(1-\mu/L)^{n}$ with the commonly proposed step size $\eta=1/L$ in discrete time based on our assumptions, Nesterov’s method achieves decay like $(1-\sqrt{\mu/L})^{n}$ with $\eta=1/L$. Since $\mu/L\leq 1$, Nesterov’s method converges much faster than gradient descent.
@@
However, if the eigenvalues of $D^{2}f$ are as negative as $-L$, the step size $\eta=\mu/L^{2}$ is so small that it does not improve upon gradient descent with step size $\eta=1/L$ since $\sqrt{\mu\eta}=\mu/L$ in this case. Thus, if $f$ is too far from being convex, acceleration may not be achievable in discrete time. Notably, especially close to the set of global minimizers, we can picture $\varepsilon$ as small compared to $L$.
3.3 Stochastic optimization in discrete time
In typical applications in deep learning, the gradient $\nabla f$ of the objective function/loss function $f$ is prohibitively expensive to evaluate, but we have access to stochastic estimates of the true gradient. In this section, in addition to the assumptions of Theorem 11, we assume that we are given a probability space $(\Omega,\mathcal{A},\mathbb{Q})$ and a measurable function $g:\mathbb{R}^{d}\times\Omega\to\mathbb{R}^{d}$ such that
${\mathbb{E}}_{\omega\sim\mathbb{Q}}\big{[}g(x,\omega)\big{]}=\nabla f(x),\qquad{\mathbb{E}}_{\omega\sim\mathbb{Q}}\big{[}\|g(x,\omega)\|^{2}\big{]}<+\infty\qquad\forall\ x\in\mathbb{R}^{d}.$ (8)
@@ 3.3 Stochastic optimization in discrete time
${\mathbb{E}}_{\omega\sim\mathbb{Q}}\big{[}\|g(x,\omega)-\nabla f(x)\|^{2}\big{]}\leq\sigma_{a}^{2}+\sigma_{m}^{2}\|\nabla f(x)\|^{2}\qquad\forall\ x\in\mathbb{R}^{d}.$ (9)
@@
Let us consider the purely additive case first. We follow the scheme (5), but we replace the deterministic gradient $\nabla f(x_{n}^{\prime})$ by $g_{n}=g(x_{n}^{\prime},\omega_{n})$ where $\omega_{0},\omega_{1},\dots$ are drawn from $\Omega$ independently of each other and the initial condition $x_{0}$ with law $\mathbb{Q}$. This framework allows e.g. for minibatch sampling (but assumes that all batches are drawn independently of each other from the dataset).
@@
Thus Nesterov’s method reduces $f(x_{n})$ below a ‘noise level’ proportional to $\sigma_{a}^{2}\sqrt{\eta/\mu}$ at a linear rate $(1-\sqrt{\mu\eta})^{n}$ in the presence of additive noise. The analogous bound for stochastic gradient descent was obtained (in the more general setting of PL functions) in (Karimi et al., 2016, Theorem 4) as
@@
With a fixed positive learning rate $\eta$, generally $f(x_{n})\not\to 0$ unless $\sigma_{a}=0$. We therefore consider a sequence of decreasing step sizes.
@@
$x_{n}^{\prime}=x_{n}+\sqrt{\alpha}v_{n},\qquad x_{n+1}=x^{\prime}_{n}-\eta g_{n},\qquad v_{n+1}=\rho_{n}\big{(}v_{n}-\sqrt{\alpha}g_{n})$ (10)
@@
[Additive and multiplicative noise] Assume that $f,P,x^{*}$ are as in Theorem 11 and that $g$ is a family of gradient estimators such that (8) and (9) hold for some $\sigma_{a},\sigma_{m}\geq 0$. Assume that the sequences $x_{n},x_{n}^{\prime},v_{n}$ are generated by the AGNES scheme (10) with parameters
$0<\eta\leq\frac{1}{L(1+\sigma_{m}^{2})},\qquad\rho=\frac{1-\sqrt{\frac{\mu\eta}{1+\sigma_{m}^{2}}}}{1+\sqrt{\frac{\mu\eta}{1+\sigma_{m}^{2}}}},\qquad\alpha=\frac{1-\sqrt{\mu(1+\sigma_{m}^{2})\eta}}{1-\sqrt{\mu(1+\sigma_{m}^{2})\eta}+\sigma_{m}^{2}}\,\eta.$
@@
The proof of Theorem 15 is given in Appendix B.4. While at a glance it appears that the multiplicative noise is helping us reduce the additive error term, this is merely a consequence of the small learning rate which is forced upon us. The condition on the negative eigenvalues is relaxed to $\varepsilon\leq\sqrt{\mu L}(1+\sigma_{m}^{2})$ with the largest admissible step size $\eta=1/(L(1+\sigma_{m}^{2}))$ as the issues stemming from tangential drift pale in comparison to those stemming from stochastic gradient estimates. For comparison, if we chose the same $\eta$ in Theorem 11, we could only allow for $\varepsilon\leq\sqrt{\mu L}\,\sqrt{1+\sigma_{m}^{2}}<\sqrt{\mu L}(1+\sigma_{m}^{2})$, but we would obtain a rate of convergence of $1-\sqrt{\mu/{L(1+\sigma_{m}^{2})}}$. In the stochastic case, we only achieve $1-\sqrt{\mu/L}/(1+\sigma_{m}^{2})$. Thus the limiting factor is the stochastic noise, not the geometry of $f$.
@@
4 Conclusion
We have proved that first order momentum-based methods accelerate convergence in a more general setting than convex optimization with many geometric features motivated by loss landscapes encountered in deep learning. The models we studied include the heavy ball ODE and deterministic and stochastic optimization schemes in discrete time under various noise assumptions.
The most cumbersome limitation of our convergence guarantees is the assumption that the closest point projection onto the set of minimizers is affine linear in the discrete time setting, i.e. the derivative $D\pi$ is constant in space. Future work will focus on relaxing this assumption.
Appendix A Proofs of Lemmas 5 and 7: Geometry of the energy landscape
@@
is convex. In particular, the derivative matrix $D\pi=D^{2}\xi$ is symmetric and positive semi-definite and thus
@@
Step 2. The geometry of $f$. Since $\mathcal{M}$ is the set of minimizers of $f$, the Hessian $D^{2}f(z)$ is positive semi-definite for every $z\in\mathcal{M}$. As $f$ is constant on $\mathcal{M}$, we for every curve $\gamma:(-\varepsilon,\varepsilon)\to\mathcal{M}$ we have
@@ Appendix B Proofs of Acceleration in Optimization
B.1 Proof of Theorems 6 and 8: Optimization in continuous time
@@
Next, using the $\mu$-strong aiming condition, we have
@@
Substituting this into the bound on ${\mathcal{L}}^{\prime}(t)$ and simplifying gives:
@@
B.2 Proof of Theorem 11: Acceleration in Discrete Time (Deterministic setting)